If you invest 50000 at an annual investment rate of 7.5% how much will your investment be worth after 11 years if your investment is compounded quarterly
Answer by DrBeeee(684)

Use the formula
(1) B = P*(1 + r/4)^(4*t) where r ii the annual rate written as a decimal and t is the number of years.
The reason for the four is due to the quarterly compounding. (If it was monthly use 12 instead of 4 in the rate and exponent.) Using values we get
(2) B = 50,000*(1 + 0.075/4)^(4*11) or
(3) B = 50,000*(1.01875)^44 or
(4) B = 50,000*(2.2645...) or
(5) B = 113,225.79
Answer: Your $50,000 investment is now worth $113,225.79